OpenStudy (aravindg):

\[\sqrt{[(2x ^{2}+x+2)(x ^{2}+3x+1)]}+2\sqrt{[(x ^{2}+3x+1)(2x ^{2}+x+2)]}=3\]

myininaya (myininaya):

\[3\sqrt{(2x^2+x+2)(x^2+3x+1)}=3\]

myininaya (myininaya):

\[\sqrt{(2x^2+x+2)(x^2+3x+1)}=1\]

myininaya (myininaya):

\[(2x^2+x+2)(x^2+3x+1)=1\]

OpenStudy (anonymous):

first of all you square both side so you end up with (2x^2+x+2)(x^2+3x+1)+4(x^2+3x+1)(2x^2+x+2)=9 and just from there you got to foil and all the good stuff..

myininaya (myininaya):

\[2x^4+6x^3+2x^2+x^3+3x^2+x+2x^2+6x+2=1\]

myininaya (myininaya):

\[2x^4+7x^3+7x^2+7x+2-1=0\]

OpenStudy (aravindg):

......................

myininaya (myininaya):

\[2x^4+7x^3+7x^2+7x+1=0\]

myininaya (myininaya):

whats wrong?

OpenStudy (aravindg):

whats the answer??

myininaya (myininaya):

solve that equation and you will find it try to see if there are any rational zeros

OpenStudy (aravindg):

thrs a problem in answer i dont know where

myininaya (myininaya):

i don't think it has any rational zeros

myininaya (myininaya):

therefore I would say use a graphing calculator in approximate the solutions but we need to check in the original equation